SpaceRef

SpaceRef


NASA Space Station On-Orbit Status 9 February 2005

Status Report From: NASA HQ/SpaceRef/NASA Watch
Posted: Wednesday, February 9, 2005

image SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by SpaceRef.com (copyright © 2005) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.  

Continuing the current outfitting of the Russian ASN-M satellite navigation system in the SM for the European ATV (Automated Transfer Vehicle), extending through 2/11, FE Salizhan Sharipov today installed two metal grounding straps between the NVM-1,2 navigation computing units mounted yesterday, and the interior frame.  The work was again supported by S-band tagup with ground specialists.   [In tomorrow s ASN activity, Salizhan will connect the ASN-M to the SUBA onboard equipment control system and on 2/11 to the SBI onboard measuring system, completing the installation.  Although not yet operational, ASN-M uses GLONASS satellites (the Russian GPS equivalent) to update the ISS state vector (SV, position & velocity plus time) without using the ground (which currently has to uplink daily SV updates) or requiring SV transfers from the U.S. segment (USOS) from time to time.  The ASN equipment was originally factory-installed in the SM but was found faulty and had to be returned to the ground.  After repair it was shipped again to the station on Progress 11 and re-installed by Yuri Malenchenko on 7/8/03, followed by various troubleshooting attempts and the current ATV mods.]

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

Both crewmembers completed the standard emergency medical CHeCS OBT (Crew Health Care Systems/on-board training) drill, a 30-min. exercise to refresh the CMO (Crew Medical Officer) s acuity in using HMS (Health Maintenance Systems) equipment, including applying ACLS (advanced cardiological life support) in an emergency.  [The proficiency drill was guided by a number of training videos and concluded with a self-assessment questionnaire, focused on re-familiarization with the RSP (respiratory support pack), defibrillator, and ALSP (advanced life support pack) components.]

Yesterday s regeneration of the two remaining expended EMU Metox (metal oxide) recyclable CO2 absorption canisters in the Quest Airlock (A/L) concluded successfully last night.  However, the temperature survey of the Metox oven and the A/L, planned in conjunction with the bakeout regeneration, was cancelled when cabin temperatures did not reach the steady state desired for the survey.   [The higher cabin temperature was the result of the ITCS (Internal Thermal Control System) being in single loop (LTL, low temperature loop), instead of dual loop.  The LTL temperature did not decrease to the pre-commanded setpoint because of the configuration of the upstream MTL TWMV (moderate temperature loop/three-way mixing valve).  Reconfigurations by the ground corrected the problem, but too late for yielding steady (relaxed) data for the planned temperature sampling.  Since all Metox canisters onboard are now regenerated for the LF-1 spacewalks, this activity will not be rescheduled for the current crew.]

As a result of the elevated temperature during Metox regen and the subsequent ground-commanded lowering of both LTL temperature and CCAA (common cabin air assembly) setpoint, some humidity condensed in the CCAA ducting, causing the air conditioner to shut down to protect the CDRA (CO2 removal assembly) filter beds.   Reactivation by the ground of CCAA and TCCS (trace contaminant control subsystem), also shut down by the protective software, was nominal.

After completion of Metox regen last night, Leroy Chiao today spent several hours in the A/L on consolidating, inventorying and stowing equipment as an important preparatory part of LF-1 prepacking.   [The activity was broken down in several tasks, with top priority being applied to gear consolidation and stowage, followed by EMU equipment bag auditing, utility bag and EMU servicing kit inventory, and EMU ORU and SPCE (service, performance, and checkout equipment) maintenance kit audit.]

In an additional two-hour task of EVA gear preparation for future use, CDR Chiao worked on three PHA (Prebreathe Hose Assembly) kits, one of them with PHA spares.   [Leroy inspected the kits contents for damage, torqued down hose fittings, covered or capped QDs (quick disconnects) for protection, and cleaned QDMs (quick don masks). Two out-of-life QDMs were stored in a return-to-Houston bag.]

With the Elektron O2 generator remaining off, Sharipov worked on the Russian harmful impurities removal system (BMP), starting the "bake-out" cycle to vacuum on absorbent bed #1 of the regenerable dual-channel filtration system.   [Regeneration of each of the two cartridges takes about 24 hours.  Elektron was deactivated on 2/7 after ppO2 (oxygen partial pressure) approached the upper concentration limit of 24.1% set by Flight Rule as a flammability safety margin.  Last evening, ppO2 was 174.5 mmHg/Torr, or 23% O2 concentration, respectively.]

CDR/SO Chiao deployed two passive FMK (formaldehyde monitoring kit) sampling assemblies in the Lab (below CEVIS) and SM (most forward handrail), to catch any atmospheric formaldehyde on a collector substrate for subsequent laboratory analysis.  (Last time done: 1/7).   [The regular periodic surface and air sampling is scheduled for tomorrow.]

The crew yesterday completed the ADUM OPE (Advanced Diagnostic Ultrasound in Microgravity/On-board Proficiency Enhancer) session, along with the scheduled HRF (Human Research Facility) laptop file management, in preparation for Salizhan s bone (Z) and abdominal (B) scans on 2/11 (Friday).   [However, the OPE file was not downlinked due to problems with the File Manager, which crashed twice.  Work is underway to get the OPE file down through OCA comm.]

For his next Saturday Science program on 2/12, Dr. Chiao has selected transferring and consolidating equipment kits for the HRF FOOT (Foot/Ground Reaction Forces during Space Flight) experiment.

The CDR conducted the routine SOZh/ECLSS servicing/inspection in the SM, including ASU toilet facilities maintenance, while the FE, working off the Russian task list, prepared the regular IMS delta file for the daily automated export/import to the three IMS databases on the ground.

Also from the discretionary job jar , Salizhan completed the regular daily inspection of the BIO-5 Rasteniya-2 ("Plants-2") experiment that researches growth and development of plants under spaceflight conditions in the Lada-5 greenhouse.

Leroy did his daily checkup of the Total Dose reading and End File values of the TEPC (Tissue Equivalent Proportional Counter), which he relocated to the Node on 1/31, and called the data down at the evening DPC (daily planning conference).   [This is currently a daily requirement since the UOP (utility outlet panel) near the TEPC s temporary location (for two weeks) does not allow automated telemetry monitoring by the ground.]

The crew worked out in accordance with their daily 2.5-hr. physical exercise protocol on TVIS treadmill (aerobic), RED exerciser (anaerobic) and VELO cycle with bungee cord load trainer (both aerobic and anaerobic).  Salizhan's daily protocol currently prescribes a four-day microcycle exercise with 1.5 hr on the TVIS (today: Day 2 of a new set) and one hour on VELO.

Chiao then transferred the daily TVIS and RED exercise data files to the MEC (medical equipment computer) for downlink, as well as the daily wristband HRM (heart rate monitor) data, followed by their erasure on the HRM storage medium.

Sharipov broke out and set up the equipment for tomorrow s scheduled Russian PZE MO-9 Urolux biochemical urine test.

After yesterday s successful attitude transition to XPOP TEA (x-axis perpendicular to orbit plane/torque equilibrium attitude) at 10:47am EST and return of control authority to the U.S. CMGs (control moment gyroscopes) at ~11:07am, an incorrect ground command caused unexpected CMG saturation (~11:26am) and loss of AC (attitude control).  AC was taken over automatically by the Russian thrusters and later returned to CMG Momentum Management (~4:54pm).   [The excursion cost an estimated 14 kg of propellants on top of the 9 kg for the original maneuver to XPOP.]

Today's CEO (crew earth observations) photo targets, limited in the current XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in ram), were Calcutta, India (this overpass provided an opportunity to capture haze over the Calcutta region and the Bay of Bengal.  Looking to the left and slightly ahead of track for potential smog/haze contrasting against the waters of the Bay.  Photography of smog and haze is useful for qualitative assessment of airflow patterns and atmospheric opacity), Po Valley, Italy (Dynamic Event.  Last week s ban on automobile use in Vicenza, Italy is over.  Looking to left of track along the southern boundary of the Alps for Vicenza [northwest of Venice].  Images of haze and smog in this region [including the Po Valley] are useful to assess the impacts, if any, of the automobile ban on regional air pollution levels), and Internal waves, E & W Florida coasts (weather was predicted to be clear over the western Florida coastline.  Looking to the right of track along the Gulf coast for internal waves; the sunglint point will be close to shore and near to the tip of the Florida peninsula).

CEO photography can be viewed and studied at the websites:

See also the website "Space Station Challenge" at:

To view the latest photos taken by the expedition 10 crew visit:

Expedition 10 Flight Crew Plans can be found at http://spaceflight.nasa.gov/station/timelines/

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

  • Upcoming Key Events:
  • ISS Reboost -- 2/15 (~8:22pm EST, ~1.8 m/s; phasing for 17P launch);
  • Progress M-51 (16P) undocking & destructive reentry -- 2/26/05;
  • Progress M-52 (17P) launch -- 2/28/05.
  • EVA-13 -- 3/25/05;
  • Soyuz TMA-6 (10S) launch -- 4/15/05 with Expedition 11 (CDR Sergei Krikalev, FE/SO John Phillips);
  • Soyuz TMA-5 (9S) undock -- 4/25/05 with Exp. 10 crew (after 193 days on orbit, 191 days on board ISS);
  • Progress M-53 (18P) launch -- 6/10/05;
  • Progress M-54 (19P) launch -- 8/24/05;
  • Soyuz TMA-7 (11S) launch -- 9/27/05.

ISS Location NOW

Full Size/Update
Real Time ISS Tracker - More Links

ISS Orbit  (as of last night, 4:50pm EST [= epoch]):

  • Mean altitude -- 357.0 km
  • Apogee height -- 363.6 km
  • Perigee height -- 350.3 km
  • Period -- 91.68 min.
  • Inclination (to Equator) -- 51.64 deg
  • Eccentricity -- 0.0009822
  • Solar Beta Angle -- -47.9 deg (magnitude decreasing)
  • Orbits per 24-hr. day -- 15.70
  • Mean altitude loss in last 24 hours -- 60 m
  • Revolutions since FGB/Zarya launch (Nov. 98) -- 35551

 

ISS Altitude History

Apogee height -- Mean Altitude -- Perigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html. In addition, information on International Space Station sighting opportunities can be found at http://spaceflight.nasa.gov/realdata/sightings/ on NASA's Human Spaceflight website. The current location of the International Space Station can be found at http://science.nasa.gov/temp/StationLoc.html at NASA's Marshall Space Flight Center. Additional satellite tracking resources can be found at http://www.spaceref.com/iss/tracking.html.

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.